8. Price Floors, Production Subsidies and Quotas Developed countries often intervene in their agricultural industries, using price floors, production subsidies or quotas (supply management). As an economist in the Department of Agriculture you have estimated the demand to be P = 300 - 4Qd and supply to be P = 100 +Qs for the chicken industry. You have been asked to evaluate three policy choices. Quantities are in tons. 8.1.To begin with, there are no interventions. Find the equilibrium P and Q. Also find the total revenue (TR) of the suppliers. 8.2. Option 1: Price floor=$160, the government buys up any surplus (excess supply). Find Qd, Qs, surplus, TR of the suppliers, and the cost to the government. 8.3. Option 2: Production subsidy Find the amount of the production subsidy per unit required such that the TR of firms is the same as under the price floor. Also find how much the consumers pay per unit and the total cost to the government. 8.4. Option 3: Quota=35 Find the corresponding consumer price and the total revenue of the suppliers at this quantity supplied. 8.5. In your capacity as an economist working for the government, which of the three options would you recommend? If you were working as an economist for the farmers, which option would you recommend? Explain.
Answer #1

Here: Deman: P = 300 - 4Qd3 - 0 Supply :P = 100 + Qs - 1 8.1 Ar equiübrium, Damand = Supply Or, 300 - 4Qd = 100 fas 0v,5Q - 28.3) her there be a production subsidy of s per unit. BTP 9600 = P Q (ITA) or, a (to) = 9697 24 evaluate at Po, Qo] apxity

8.5) If I was an economist working for the government, my objectives would be:

  • To minimize the cost to government, and
  • To maximise welfare.

Thus, I would have opted for the subsidy, as in this case the cost is minimum (100 < 4000) and also consumers enjoy their old prices.

If I was an economist working for the farmers, my objectives would be to maximise farmers’ income. This occurs when at the price floor of $160. Thus, I would go ahead for the price floor.